# PPC-GPT: Federated Task-Specific Compression of Large Language Models via Pruning and Chain-of-Thought Distillation

### v1

@@ 3.3 Synthetic Data Generation
    Question Generation. We prompt $LLM_{syn}$ to create a new question, starting from a perturbed question. To enhance the validity of these new created questions, we enforce three guidelines within the prompt: (1) the new question needs to conform to common knowledge, (2) it must be solvable on its own, independent of the original question, and (3) it should not contain any answer responses. Furthermore, we establish specific formatting standards for both questions and answers, customized to suit the needs of various datasets Li et al. (2024).
@@
Detailed prompt designs are presented in Appendix B. The generated synthetic data and their rationales are then employed for model pruning and retraining on the server-side.
3.4 Layer-Wise Structured Pruning
In neural networks exhibiting high redundancy, it is plausible that certain layers contribute minimally to the ultimate performance of the model. This phenomenon can be attributed to the homogeneous functionalities of these layers relative to others within the network architecture. LLMs are no exception, with varying levels of redundancy observed across their layers, particularly more pronounced in deeper layers. To eliminate these redundant layers, we require a metric that is inherent to the model itself for assessing the significance of each layer. This metric should be capable of quantifying the essentiality of each layer in relation to the model’s overall functionality and performance.
To quantify the impact of each layer, we use a novel metric termed "Block Influence"(BI), which is proposed in the ShortGPT Men et al. (2024). This metric is grounded in the hypothesis that a transformer block’s significance is directly proportional to the extent it modifies the hidden states. Mathematically, the BI score for the $i^{th}$ block is computed as:
$\text{BI}_{i}=1-\mathbb{E}_{X,t}\left[\frac{X_{i,t}^{T}X_{i+1,t}}{||X_{i,t}||_{2}||X_{i+1,t}||_{2}}\right],$ (6)
@@ 3.4 Layer-Wise Structured Pruning
The original BI method Men et al. (2024) relies solely on input and task label information, processed through a single forward pass: $f_{\theta_{o}}(x_{i})\rightarrow y_{i}$. We further extend the BI computation to encompass two distinct facets of influence: $f_{\theta_{o}}(x_{i})\rightarrow y_{i}$ and $f_{\theta_{o}}(x_{i})\rightarrow r_{i}$. This enhancement not only facilitates the prediction of task labels but also enables the generation of corresponding rationales based on the inputs. Our novel BI score is determined as follows:
@@
A higher BI score indicates a more crucial layer. An example is shown in the Figure 1(b). We arrange the layers in ascending order based on their BI scores and proceed to eliminate those with the lower BI scores.
@@ 3.5 Retraining
Server-side Retraining. On the server side, we utilize the synthetic dataset $\mathcal{D}_{s}$, as described in Section 3.3, to retrain the pruned model, $SLM_{t}$. We propose COT knowledge distillation, guided by rationales generated by $LLM_{syn}$, to enhance the performance of $SLM_{t}$. Formally, we conceptualize the learning process with rationales as a multi-task learning problem Zhang and Yang (2021); Wei et al. (2022); Hsieh et al. (2023). Specifically, we train the model $f_{\theta_{t}}(x_{i})\rightarrow(y_{i},r_{i})$ to achieve not only the prediction of task labels but also the generation of corresponding rationales based on textual inputs. This multi-task training ensures that our model produces not only accurate predictions but also insightful justifications for its decisions, thereby enhancing the model’s transparency and explainability. The multi-task learning objective can be formulated as follows:
@@
$\mathcal{L}_{\text{Label}}(\theta_{t};\mathcal{D}_{s})=\mathbb{E}_{(x,y)\sim\mathcal{D}_{s}}\ell_{\text{CE}}(f_{\theta_{t}}(x),y)$ (9)
@@
$\mathcal{L}_{\text{Rationale}}(\theta_{t};\mathcal{D}_{s})=\mathbb{E}_{(x,r)\sim\mathcal{D}_{s}}\ell_{\text{CE}}(f_{\theta_{t}}(x),r)$ (10)
where $\ell_{\text{CE}}$ denotes the cross-entropy loss, $f_{\theta_{t}}(\cdot)$ denotes the $SLM_{t}$ model.
Client-side Retraining. On the client side, we utilize local private data $\mathcal{D}$ to further retrain the pruned model, $SLM_{t}$, once it has been received from the server. Our work encompasses conventional training, leveraging ground truth labels to further enhance the performance of $SLM_{t}$. Formally, the label prediction loss for this dataset $\mathcal{D}$ is formulated as follows:
$\mathcal{L}_{\text{Label}}(\theta_{t};\mathcal{D})=\mathbb{E}_{(x,y)\sim\mathcal{D}}\ell_{\text{CE}}(f_{\theta_{t}}(x),y)$ (11)
@@ 4.1 Setup
We have devised a scenario to assess the performance of the PPC-GPT framework across various text generation tasks. This setup employs a client-server architecture, where the server hosts a LLM for synthetic data generation, denoted as $LLM_{syn}$. Specifically, we have selected LLaMa3-70B Dubey et al. (2024) for this purpose. For model pruning, we utilize LLaMA2-7B Touvron et al. (2023b) and OPT-6.7B Zhang et al. (2022) as the source models, denoted as $LLM_{o}$. In the default setting, the privacy budget $\epsilon=3$, and the synthetic data ratio is 8.
Datasets and Evaluation Metrics. We conduct a comparative evaluation of PPC-GPT on QA datasets. Specifically, we include CommonsenseQA (CQA) Talmor et al. (2018), OpenBookQA (OBQA) Mihaylov et al. (2018), ARC-C Clark et al. (2018), ARC-E Clark et al. (2018), FiQA-SA Maia et al. (2018). For these datasets, we primarily use Accuracy as the evaluation metric. It’s worth noting that in our experiments, all methods undergo zero-shot evaluation and we use the lm-evaluation-harness package Gao et al. (2023).
@@
  • DP-Instruct-C Yu et al. (2024), where the client finetunes generator (e.g., LLaMa2-1.3B) with DP-SGD and using synthetic datasets generated from generator to prune $LLM_{o}$ and subsequently fine-tunes the pruned model with the private dataset.
@@ 4.2 Main Results
In our experiments, we extensively evaluated the performance of the proposed PPC-GPT framework across various text generation tasks. Notably, given that current structured pruning methods typically reduce parameters by no more than 30%, we conducted experiments with approximately 30% of the parameters pruned. Additional experiments exploring different parameter reduction proportions will be discussed in Section 5.5.
@@
5 Ablation Study
5.1 Impact of Different Privacy Budgets
@@
Table 2: Comparison of PPC-GPT’s performance across different privacy budgets $\epsilon$. S denotes the performance of target $SLM_{t}$ on the server-side, while C represents the performance of target $SLM_{t}$ on the client-side.
5.2 Impact of Different Synthetic Data
@@
5.3 Impact of Server-Side Retraining
@@
5.4 Impact of Different Importance Metric
@@
5.5 Impact of Different Model Pruning Ratio
@@ 6 Conclusions
In this study, we introduce PPC-GPT, a federated framework designed to compress LLMs into task-specific SLMs while preserving privacy. This is achieved through pruning and COT distillation. PPC-GPT employs a server-client architecture, wherein client data perturbed with DP is transmitted to the server for the generation of synthetic data. This synthetic data, along with its associated rationales, is subsequently used for both the pruning of LLMs and the retraining of pruned SLMs. Our experimental findings show that PPC-GPT successfully compresses LLMs into highly competitive SLMs, with a strong emphasis on safeguarding data privacy.
@@
While PPC-GPT shows promising results in compressing LLMs into task-specific SLMs while ensuring data privacy, it has several limitations. Firstly, PPC-GPT relies on the LLM’s robust chain-of-thought capabilities to generate high-quality synthetic data and rationales. If the LLM lacks such capabilities, the performance of the compressed SLMs may be impacted. Furthermore, as observed in our experiments, the performance of PPC-GPT tends to degrade with higher pruning ratios. This indicates that optimizing the pruning strategy to strike a better balance between model size and performance remains an open challenge.
Appendix A Implementation Details
A.1 Hyperparameter Settings
@@
A.2 Data Splitting
For the datasets, all splits (training, validation, and test) were downloaded from HuggingFace Lhoest et al. (2021).
A.3 Dataset Licenses
All the datasets were downloaded from HuggingFaceLhoest et al. (2021) and under Apache License, Version 2.0.
A.4 Machine Configuration
@@
Appendix B Synthetic Prompt Templates

### v2

@@ 4.2 CoT-guided Synthetic Data Generation
    Question Generation. We prompt $LLM_{syn}$ to create a new question, starting from a perturbed question. To enhance the validity of these new created questions, we enforce three guidelines within the prompt: (1) the new question needs to conform to common knowledge, (2) it must be solvable on its own, independent of the original question, and (3) it should not contain any answer responses. Furthermore, we establish specific formatting standards for both questions and answers, customized to suit the needs of various datasets li2024common.
@@
Detailed prompt designs are presented in Appendix C. The generated synthetic data and their rationales are then employed for model pruning and retraining on the server-side.
Figure 2: CoT-Guided Synthetic Data Generation.
4.3 Rationale-Aware Structured Pruning
LLMs exhibit layer-wise redundancy, with deeper layers often showing higher levels of functional overlap. To identify and remove redundant layers effectively, we need a quantitative metric that can assess each layer’s contribution to the model’s performance. This intrinsic metric should evaluate both the layer’s individual importance and its interaction with other layers in maintaining the model’s overall functionality.
To quantify the impact of each layer, we use a novel metric termed "Block Influence" (BI), which is proposed in the ShortGPT men2024shortgpt. This metric is grounded in the hypothesis that a transformer block’s significance is directly proportional to the extent it modifies the hidden states. Mathematically, the BI score for the $i^{th}$ block is computed as:
[…]
@@ 4.3 Rationale-Aware Structured Pruning
The original BI method men2024shortgpt relies solely on input and task label information, processed through a single forward pass: $f_{\theta}(x^{s}_{i})\rightarrow y^{s}_{i}$. We further extend the BI computation to encompass two distinct facets of influence: $f_{\theta}(x^{s}_{i})\rightarrow y^{s}_{i}$ and $f_{\theta}(x^{s}_{i})\rightarrow r^{s}_{i}$. This enhancement not only facilitates the prediction of task labels but also enables the generation of corresponding rationales based on the inputs. Our novel BI score is determined as follows:
@@
A higher BI score indicates greater layer importance in the model architecture. As illustrated in Figure 3, we leverage these BI scores to guide our pruning strategy: layers are arranged in ascending order based on their BI scores, and those with lower scores are systematically removed to obtain the pruned model structure SLM $f_{\phi}$.
Figure 3: Layer Importance Example: The significance of each layer, as indicated by the BI (Block Influence) value of LLaMA2-7B on the OBQA dataset, based on the PPC-GPT framework.
@@ 4.4 Retraining
Server-side Retraining. On the server side, we utilize the synthetic dataset $\mathcal{D}_{s}$, as described in Section 4.2, to retrain the pruned model SLM $f_{\phi}$. We propose CoT knowledge distillation, guided by rationales generated by $LLM_{syn}$, to enhance the performance of SLM $f_{\phi}$. Formally, we conceptualize the learning process with rationales as a multi-task learning problem zhang2021survey; wei2022chain; hsieh2023distilling. Specifically, we train the model $f_{\phi}(x^{s}_{i})\rightarrow(y^{s}_{i},r^{s}_{i})$ to achieve not only the prediction of task labels but also the generation of corresponding rationales based on textual inputs. This multi-task training ensures that our model produces not only accurate predictions but also insightful justifications for its decisions, thereby enhancing the model’s transparency and explainability. The multi-task learning objective can be formulated as follows:
@@
$\mathcal{L}_{\text{Label}}(\phi;\mathcal{D}_{s})=\mathbb{E}_{(x^{s},y^{s})\sim\mathcal{D}_{s}}\ell_{\text{CE}}(f_{\phi}(x^{s}),y^{s})$ (9)
@@
$\mathcal{L}_{\text{Rationale}}(\phi;\mathcal{D}_{s})=\mathbb{E}_{(x^{s},r^{s})\sim\mathcal{D}_{s}}\ell_{\text{CE}}(f_{\phi}(x^{s}),r^{s})$ (10)
where $\ell_{\text{CE}}$ denotes the cross-entropy loss.
Client-side Retraining. On the client side, we utilize local private data $\mathcal{D}$ to further retrain the pruned model, SLM $f_{\phi}$, once it has been received from the server. Our work encompasses conventional training, leveraging ground truth labels to further enhance the performance of SLM $f_{\phi}$. Formally, the label prediction loss for this dataset $\mathcal{D}$ is formulated as follows:
$\mathcal{L}_{\text{Label}}(\phi;\mathcal{D})=\mathbb{E}_{(x,y)\sim\mathcal{D}}\ell_{\text{CE}}(f_{\phi}(x),y)$ (11)
@@ 5.1 Setup
We have devised a scenario to assess the performance of the PPC-GPT framework across various text generation tasks. This setup employs a client-server architecture, where the server hosts an auxiliary LLM for synthetic data generation, denoted as $LLM_{syn}$. Specifically, we have selected LLaMA3-70B dubey2024llama for this purpose. For model pruning, we utilize LLaMA2-7B touvron2023llama2 and OPT-6.7B zhang2022opt as the source models, denoted as $LLM_{o}$. In the default setting, the privacy budget $\epsilon=3$, and the synthetic data ratio is 8.
Datasets and Evaluation Metrics. We conduct a comparative evaluation of PPC-GPT on QA datasets. Specifically, we include CommonsenseQA (CQA) talmor2019commonsenseqa, OpenBookQA (OBQA) mihaylov2018can, ARC-C clark2018think, ARC-E clark2018think, FiQA-SA maia201818. For these datasets, we primarily use Accuracy as the evaluation metric. It’s worth noting that in our experiments, all methods undergo zero-shot evaluation and we use the lm-evaluation-harness package eval-harness.
@@
  • DP-Instruct-C yu2024privacy, where the client finetunes generator (e.g., LLaMA2-1.3B) with DP-SGD and using synthetic datasets generated from generator to prune $LLM_{o}$ and subsequently fine-tunes the pruned model with the private dataset.
@@ 5.2 Main Results
In our experiments, we extensively evaluated the performance of the proposed PPC-GPT framework across various text generation tasks. Notably, given that current structured pruning methods typically reduce parameters by no more than 30%, we conducted experiments with approximately 30% of the parameters pruned. Additional experiments exploring different parameter reduction proportions will be discussed in Section 5.3.5.
@@
5.3 Ablation Study
5.3.1 Impact of Different Privacy Budgets
@@
Table 2: Comparison of PPC-GPT’s performance across different privacy budgets $\epsilon$. S denotes the performance of target SLM on the server-side, while C represents the performance of target SLM on the client-side.
5.3.2 Impact of Different Synthetic Data
@@
5.3.3 Impact of Server-Side Retraining
@@
5.3.4 Impact of Different Importance Metric
@@
5.3.5 Impact of Different Model Pruning Ratio
@@ 6 Conclusions
In this paper, we introduced PPC-GPT, a novel federated framework for compressing LLMs into task-specific SLMs while preserving privacy. Our framework integrates four key components: exponential mechanism-based data perturbation, CoT-guided synthetic data generation, rationale-aware structured pruning, and CoT-based knowledge distillation. Experiments demonstrate that PPC-GPT effectively compresses LLMs while maintaining comparable performance and ensuring privacy protection. This work provides a practical solution for deploying LLMs in resource-constrained, privacy-sensitive scenarios.
@@
While PPC-GPT shows promising results in compressing LLMs into task-specific SLMs while ensuring data privacy, it has several limitations. Firstly, PPC-GPT relies on an auxiliary LLM with robust CoT capabilities to generate high-quality synthetic data and rationales. These synthetic data are crucial for guiding the structured pruning and retraining processes of both the source LLM (the model slated for compression) and the target SLM (the compressed model). If the auxiliary LLM lacks sophisticated CoT reasoning abilities, the quality and diversity of the generated synthetic data may be compromised, which in turn could adversely affect the performance of the compressed SLMs. This limitation underscores the importance of selecting or pre-training an auxiliary LLM with strong CoT capabilities when deploying PPC-GPT. However, it’s important to note that the source LLM (the model slated for compression) does not necessarily require CoT capabilities. Furthermore, as observed in our experiments, the performance of PPC-GPT tends to degrade with higher pruning ratios. This indicates that optimizing the pruning strategy to strike a better balance between model size and performance remains an open challenge.
Appendix A Privacy Analysis of PPC-GPT
Our privacy protection strategy in PPC-GPT is grounded in rigorous theoretical foundations and validated through comprehensive empirical studies. The framework implements a theoretically-sound differential privacy (DP) mechanism that operates at the token-level feature space, completely eliminating the need for raw data transmission. Specifically, we adopt the exponential mechanism, which provides formal $\epsilon$-DP guarantees and has been extensively analyzed in privacy-preserving NLP literature yue2021differential; chen-etal-2023-customized; tong2025inferdpt. The theoretical privacy guarantees of this mechanism are well-established, allowing us to focus on its practical implementation and performance optimization rather than re-establishing its privacy properties.
Appendix B Implementation Details
B.1 Hyperparameter Settings
@@
B.2 Data Splitting
For the datasets, all splits (training, validation, and test) were downloaded from HuggingFace lhoest-etal-2021-datasets.
B.3 Dataset Licenses
All the datasets were downloaded from HuggingFacelhoest-etal-2021-datasets and under Apache License, Version 2.0.
B.4 Machine Configuration
@@
Appendix C Synthetic Prompt Templates